OpenStudy (anonymous):

4/9 + - 1/8 = 3/17 5/17 41/72 23/72

OpenStudy (ash2326):

@HintonCovington We have \[\frac 49 + -\frac 18\] that can be written as \[\frac 49+\frac 18\] Whenever we need to add or subtract two fraction, first step is to ensure that the two have the same denominator, here the denominators are not same, so we need to do that. Do you know LCM?

OpenStudy (anonymous):

No What Is That?

OpenStudy (anonymous):

No ash it should be written as -

OpenStudy (ash2326):

@HintonCovington I gotta bounce, my friends will help you @lgbasallote @ParthKohli

OpenStudy (anonymous):

ight

OpenStudy (ash2326):

@rebeccaskell94

OpenStudy (lgbasallote):

@ParthKohli please help here

razor99 (razor99):

last one is the correct answer

OpenStudy (anonymous):